# Simulating many-engine spacecraft: Exceeding 1 quadrillion degrees of freedom via information geometric regularization

### v3

@@ VII-C Energy efficiency
The IGR method significantly reduces energy consumed compared to the current state of the art on all platforms. The most significant reduction comes from the improved time to solution, and to lowest order, the Frontier and El Capitan energy improvements match the grind time speedups in table˜III. Due to lower power draw for WENO on Alps, we observe energy savings beyond those due to grind time speedup.
The largest improvement is realized on Frontier with a $5.38$-times improvement in energy consumed by the GPU. As described in section˜VI-C, the Frontier measurements include only GPU and GPU HBM energy and do not represent changes that may have impacted energy consumption of the CPU or CPU memory.
VIII Implications
@@
The combination of IGR, the resulting simplified algorithm, and their optimized implementation improves the efficiency of simulating compressible flow, which we measure on key metrics: time to solution, memory footprint, and energy to solution. We take advantage of the closely coupled architectures on modern supercomputing platforms, including the MI250X, MI300A, and GH200, which serve as proof of concept for efficient use of unified memory addressing.
Furthermore, our simplified algorithm is amenable to mixed-precision computation. The resulting computational savings enable the first CFD simulations with more than 200 trillion grid points and 1 quadrillion degrees of freedom, exceeding the previous largest such simulations by a factor of 20. At the same time, compared to the baseline, the improved grind times and near-ideal strong scaling of our approach enable flagship supercomputers to achieve orders of magnitude reductions of time to solution on smaller problems. This unlocks new opportunities for the integration of simulation into design optimization. Beyond flagship supercomputing, the drastic increase in grid points per device extends the capabilities of resource-constrained users.
This work focuses on a set of thruster plumes. However, the methodology shown is suited to general multiphase and multicomponent flows, spanning fields from biomedical treatments to marine and aircraft. Our work thus has the potential to aid the understanding of a broad range of physical phenomena. IGR is agnostic to the numerical discretization, as it regularizes the momentum balance equation of the associated PDE. Thus, one could apply the same strategy to other techniques, such as discontinuous Galerkin or finite difference methods.

### v4

@@ 7.3. Energy efficiency
8. Implications
@@
The combination of IGR, the resulting simplified algorithm, and its optimized implementation improves the efficiency of simulating compressible flow, as measured by key metrics: time to solution, memory footprint, and energy to solution. We leverage the closely coupled architectures of modern supercomputing platforms, including the MI250X, MI300A, and GH200, which serve as proof of concept for the efficient use of unified memory addressing.
Furthermore, our simplified algorithm is amenable to mixed-precision computation. The resulting computational savings enable the first CFD simulations with more than 200 trillion grid points and 1 quadrillion degrees of freedom, exceeding the previous largest such simulations by a factor of 20. At the same time, compared to the baseline, the improved grind times and near-ideal strong scaling of our approach enable flagship supercomputers to achieve orders of magnitude reductions of time to solution on smaller problems. This unlocks new opportunities for integrating simulation into design optimization. Beyond flagship supercomputing, the drastic increase in grid points per device extends the capabilities of resource-constrained users.
This work focuses on a set of thruster plumes. However, the methodology shown is suited to general multiphase and multicomponent flows, spanning fields ranging from biomedical treatments to marine and aviation applications. Our work thus has the potential to aid the understanding of a broad range of physical phenomena. IGR is agnostic to the numerical discretization, as it regularizes the momentum balance equation of the associated PDE. Thus, one could apply the same strategy to other techniques, such as discontinuous Galerkin or finite difference methods.